The attendance at a rugby game is as follows
Number of men : number of women : number of children = 17:9:12
Given that there are 3512 more men than women, work out the number of children attending the game.

Answer by ikleyn(52802)

.
17x men;  9x women;  and 12x children.


17x - 9x = 3512


8x = 3512


x = 3512/8 = 439


12x = 12*439 = 5268  is the number of children.     ANSWER